# Random Forests as Statistical Procedures: Design, Variance, and Dependence

### v2

@@ 3.3 Decomposing the within-resample variance term
We now refine the first term in (1), corresponding to variability induced by random splitting decisions after the resampling pattern is fixed. Define
$V_{\mathrm{in}}(x)=\mathbb{E}\!\left[\operatorname{Var}\!\left(T_{\theta}(x)\mid\mathcal{I}_{\theta},X\right)\;\middle|\;X\right].$
Conceptually, $V_{\mathrm{in}}(x)$ measures how sensitive the prediction at $x$ is to the particular way the tree is grown, holding fixed which observations are available to it. Even with the same resampled data, different choices of candidate split variables, cutpoints, or tie-breaking can route $x$ into different terminal nodes and therefore group it with different subsets of observations. Under a fixed resampling pattern, the remaining tree-level randomization affects $T_{\theta}(x)$ only through the induced averaging set $A_{\theta}(x)$. Conditioning on $A_{\theta}(x)$ therefore fixes the local grouping used to form the prediction at $x$. Applying the law of total variance conditional on $X$, with $A_{\theta}(x)$ as the intermediate conditioning object, yields
@@
The first term in (2) does not represent additional algorithmic randomness. Once both the resampling pattern and the averaging set at $x$ are fixed, the remaining tree-level randomization has been removed; variability in this component arises only from outcome randomness under the fixed-design model, propagated through the local averaging rule determined by $(\mathcal{I}_{\theta},A_{\theta}(x),X)$. Conditional on $(\mathcal{I}_{\theta},A_{\theta}(x),X)$, the weights $W_{i}(x;\theta)$ are deterministic, and the tree prediction is a fixed linear combination of outcomes. Therefore
$\operatorname{Var}\!\left(T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x),X\right)=\sum_{i=1}^{n}\sum_{j=1}^{n}W_{i}(x;\theta)\,W_{j}(x;\theta)\,\operatorname{Cov}(Y_{i},Y_{j}\mid A_{\theta}(x),X),$
which under conditional independence of outcomes across observational units reduces to
@@
Because $\mathcal{I}_{\theta}(x)$ is a component of the algorithmic randomization drawn independently of the outcome-generating process, it carries no information about $Y_{i}$ beyond $X$ and can be removed from the conditioning. The averaging set $A_{\theta}(x)$, by contrast, is determined jointly by the algorithmic randomization and the realized outcomes through data-adaptive split selection, and therefore remains in the conditioning. This is the variance of a weighted average of conditionally independent outcomes at a realized set of design points, with weights fixed by the tree structure. Taking expectations over the design yields
@@
This term captures the intrinsic statistical variability of the averaging rule: how outcome noise at the contributing observations propagates through a fixed local average. The conditional variances $\operatorname{Var}(Y_{i}\mid A_{\theta}(x),X)$ are properties of the outcome model at the realized design points, not parameters of the forest design. They enter as scaling factors that govern how much outcome dispersion is available for the weights to transmit. The design controls the weights; the outcome model controls the noise; this term measures their interaction at a realized grouping.
The second term in (2) captures variability arising purely from randomness in the induced averaging set. Even with the resampling pattern fixed, different realizations of the remaining tree-level randomization can induce different sets $A_{\theta}(x)$ and hence different local groupings of observations around $x$. Because
$\mathbb{E}[T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x),X]=\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i},$
where the weights are deterministic functions of the averaging set, this component measures variability in the prediction at $x$ induced by randomness in the grouping itself:
@@
It is zero if and only if the averaging set $A_{\theta}(x)$ is almost surely constant given $(\mathcal{I}_{\theta},X)$. This term therefore reflects instability in how the tree-growing procedure defines local neighborhoods around $x$: different partitions lead to different groupings and hence to different predictions, even when the available data are the same.
Together, the two components describe how unstable a single tree’s prediction at $x$ can be even when the same resampled data are used: the first term captures intrinsic noise propagation through a fixed averaging rule, while the second captures sensitivity to how that rule is selected.
@@ 3.4 Decomposing the resampling component
We next interpret the second term in (1),
$V_{\mathrm{out}}(x):=\operatorname{Var}\!\left(\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta},X\right]\;\middle|\;X\right),$
which captures variability induced by the resampling design across trees, conditional on the realized covariate configuration $X$. This component measures how the average prediction at $x$ changes across resampling realizations, before accounting for any additional variability from the remaining tree-level randomization.
Conditioning on the resampling indicators and using the weighted representation of the tree predictor,
@@
where $\bar{W}_{i}(x;\mathcal{I}_{\theta})$ is the expected weight assigned to observation $i$ at $x$ under the remaining tree-level randomization, given the resampling realization and the fixed covariate configuration. With this notation,
@@
Conceptually, $V_{\mathrm{out}}(x)$ isolates the effect of resampling on prediction under the design-based distribution conditional on $X$. Different resampling realizations $\mathcal{I}_{\theta}$ change which observations are exposed to the tree, which in turn alters the distribution of the induced averaging set $A_{\theta}(x)$ and hence shifts the expected local averaging rule $\bar{W}(x;\mathcal{I}_{\theta})$. Unlike $V_{\mathrm{in}}(x)$, which reflects instability in how a given resampled dataset is locally partitioned, $V_{\mathrm{out}}(x)$ reflects variability arising from changes in which observations are even eligible to contribute to the prediction at $x$.
@@ 4.1 Law of Total Covariance Decomposition
We now examine the covariance term $C_{T}(x)$ appearing in Theorem 1. From the finite-sample variance identity, $C_{T}(x)$ is the limiting contribution to predictive variability as $B\to\infty$, representing irreducible dependence between tree-level prediction rules induced by the forest design at fixed sample size. Although the trees indexed by $\theta$ and $\theta^{\prime}$ are generated independently, their predictions at $x$ are generally dependent because both act on the same realized outcomes under the fixed covariate configuration $X$ through randomized grouping mechanisms. To isolate the sources of this dependence, we condition on the resampling indicators $\mathcal{I}_{\theta}$ and $\mathcal{I}_{\theta^{\prime}}$, which determine which observations are exposed to each tree. Conditional on these indicators and on $X$, any remaining dependence reflects how the forest construction induces related prediction rules at $x$ under independent tree generation.
Applying the law of total covariance conditional on $X$ with respect to the resampling indicators yields
@@
Because the two trees are generated independently, $\mathcal{I}_{\theta^{\prime}}$ carries no information about $T_{\theta}(x)$ beyond $\mathcal{I}_{\theta}$ once $X$ is fixed, and likewise with roles reversed; we therefore suppress redundant conditioning arguments implied by independence.
The two terms in (3) correspond to distinct design mechanisms. The first term captures dependence arising from joint inclusion of observations in the averaging sets $A_{\theta}(x)$ and $A_{\theta^{\prime}}(x)$, whereby the same outcomes receive positive weight in both trees. The second term captures dependence arising from alignment of the induced prediction rules at $x$: even when the averaging sets are disjoint, independently generated trees may induce similar local averaging behavior after conditioning on their respective resampling realizations and the fixed covariate configuration.
This decomposition is exact at fixed sample size and fixed $X$. All expectations and covariances are taken with respect to the joint randomness of $(Y\mid X)$ and the tree-generating mechanism $\theta$. Resampling schemes govern the first term through joint inclusion probabilities in the averaging sets, while feature-level randomization and split selection govern the second term through stability or alignment of the induced grouping mechanism around $x$. The following subsections analyze these components in turn.
@@ 4.2 Covariance Induced by Shared Training Observations
We first analyze the leading term, $\mathbb{E}\!\left[\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X\right)\;\middle|\;X\right]$, which captures dependence arising from reuse of the same training observations in the predictions at $x$. This mechanism is present whenever the same observation receives nonzero weight in both trees at $x$. Joint inclusion in the resampling realizations is necessary but not sufficient; the observation must also lie in the terminal region containing $x$ in both trees.
Using the weighted representation $T_{\theta}(x)=\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i}$ and conditioning on $(\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X)$, the covariance expands as
$\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\;\middle|\;\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X\right)$ $=\operatorname{Cov}\!\left(\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i},\;\sum_{j=1}^{n}W_{j}(x;\theta^{\prime})\,Y_{j}\;\middle|\;\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X\right)$
$=\sum_{i=1}^{n}\sum_{j=1}^{n}W_{i}(x;\theta)\,W_{j}(x;\theta^{\prime})\,\operatorname{Cov}(Y_{i},Y_{j}\mid\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X).$
Because the resampling indicators are drawn independently of the outcome-generating process, conditioning on $(\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}})$ does not alter the joint distribution of $(Y_{i},Y_{j})$ given $X$. Moreover, under conditional independence of outcomes across observational units given the fixed covariate configuration, $\operatorname{Cov}(Y_{i},Y_{j}\mid X)=0$ for $i\neq j$, so the double sum reduces to
@@
where we define
$\sigma_{i}^{2}:=\operatorname{Var}(Y_{i}\mid X),$
the conditional outcome variance at the realized design point $X_{i}$. This quantity is a property of the fixed-design outcome model and does not depend on the forest construction. Taking expectations over the tree-generating randomization conditional on $X$ yields
@@
The joint weight factor can be decomposed as
@@
separating joint inclusion under the resampling scheme from the conditional allocation of weight induced by the local partitions. This contribution reflects dependence induced purely by repeated reuse of the same observations in the predictions at $x$. It persists regardless of whether the induced partitions agree across trees, and is identically zero only when the resampling design assigns zero probability to joint inclusion of any observation across trees.
@@ 4.3 Covariance Induced by Partition Alignment
We now analyze the second term in (3),
$\operatorname{Cov}\!\left(\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta},X\right],\mathbb{E}\!\left[T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta^{\prime}},X\right]\;\middle|\;X\right),$
which captures dependence arising from alignment of the induced local geometry around $x$ across independently generated trees. This mechanism does not require any observation to receive positive weight in both trees at $x$ in a given realization. Instead, it arises when distinct trees, potentially trained on disjoint exposed subsets, induce the same (or very similar) terminal region containing $x$ and therefore apply similar local averaging rules. Concretely, two trees trained on different observations may split on the same variables in approximately the same order at similar thresholds in a neighborhood of $x$; in that case $x$ is routed into essentially the same geometric terminal node, and each tree averages outcomes from the same covariate-defined subpopulation even if the contributing training observations are disjoint.
To formalize this, let $R_{\theta}(x)\subset\mathcal{X}$ denote the terminal region containing $x$ under tree draw $\theta$. Because only observations exposed to the tree can be routed, the averaging set defined in Section 2.2 can be written equivalently as
@@
Thus, $T_{\theta}(x)$ is a local average over exposed outcomes in the covariate-defined region $R_{\theta}(x)$. When independent trees induce similar regions $R_{\theta}(x)$ and $R_{\theta^{\prime}}(x)$, they average different observations drawn from essentially the same local subpopulation determined by those split conditions, producing dependence through aligned local geometry rather than shared outcomes.
Conditional on the resampling indicators and fixed $X$, the remaining randomness in $T_{\theta}(x)$ arises from feature subsampling and split selection, which determine the realized region $R_{\theta}(x)$ and hence the induced averaging set. Averaging over this remaining randomness induces the resampling-conditional prediction rule
$Z_{\theta}(x):=\mathbb{E}[T_{\theta}(x)\mid\mathcal{I}_{\theta},X],\qquad Z_{\theta^{\prime}}(x)=\mathbb{E}[T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta^{\prime}},X].$
Define the resampling-conditional mean weights
$\bar{W}_{i}(x;\mathcal{I}_{\theta}):=\mathbb{E}\!\left[W_{i}(x;\theta)\;\middle|\;\mathcal{I}_{\theta},X\right],\qquad i=1,\ldots,n,$
so that
$Z_{\theta}(x)=\sum_{i=1}^{n}\bar{W}_{i}(x;\mathcal{I}_{\theta})\,Y_{i},\qquad Z_{\theta^{\prime}}(x)=\sum_{i=1}^{n}\bar{W}_{i}(x;\mathcal{I}_{\theta^{\prime}})\,Y_{i}.$
The mean weights $\bar{W}_{i}(x;\mathcal{I}_{\theta})$ do not correspond to membership in a single realized terminal region. Rather, they average over the distribution of terminal regions induced by split-level randomization, assigning observation $i$ mass proportional to how frequently it is grouped with $x$ across possible tree realizations. This resampling-conditional mean weight is an intermediate quantity between the realized tree-level weight $W_{i}(x;\theta)$ and the forest-level expected weight $\mathbb{E}[W_{i}(x;\theta)\mid X]$; its introduction makes the alignment mechanism formally identifiable as a distinct component of the covariance structure. In this sense, $\bar{W}_{i}(x;\mathcal{I}_{\theta})$ represents the resampling-conditional local neighborhood around $x$ induced by the procedure.
The alignment component of the covariance is therefore
$C_{\mathrm{align}}(x):=\operatorname{Cov}\!\left(Z_{\theta}(x),Z_{\theta^{\prime}}(x)\;\middle|\;X\right),$
where the covariance is taken under the joint randomness of $(Y\mid X)$ and the independent resampling indicators $(\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}})$. Expanding yields
$C_{\mathrm{align}}(x)=\sum_{i=1}^{n}\sum_{j=1}^{n}\operatorname{Cov}(Y_{i},Y_{j}\mid X)\,\mathbb{E}\!\left[\bar{W}_{i}(x;\mathcal{I}_{\theta})\,\bar{W}_{j}(x;\mathcal{I}_{\theta^{\prime}})\;\middle|\;X\right].$
Under conditional independence of outcomes across observational units given $X$, $\operatorname{Cov}(Y_{i},Y_{j}\mid X)=0$ for $i\neq j$, and therefore
@@
where $\sigma_{i}^{2}=\operatorname{Var}(Y_{i}\mid X)$.
Expression (4) makes the alignment mechanism explicit in design-based terms. The joint expectation $\mathbb{E}[\bar{W}_{i}(x;\mathcal{I}_{\theta})\,\bar{W}_{i}(x;\mathcal{I}_{\theta^{\prime}})\mid X]$ measures whether independent trees tend, on average, to place predictive mass near $x$ on the same local subset of design points after averaging over split-level randomization. When routing near $x$ is stable across tree realizations—so that similar split conditions recur and $x$ repeatedly falls into essentially the same terminal region—the procedure concentrates on a restricted family of local regions in covariate space. The observations populating those regions, though potentially disjoint across trees, are drawn from the same conditional subpopulation defined by the shared partition geometry. Both tree predictions therefore estimate the same conditional mean, yielding positive covariance even in the absence of shared training observations.
@@
Figure 2: Two distinct design-induced dependence mechanisms at a fixed prediction point $x$. (A) Observation overlap: independently generated trees reuse the same outcome (here, $Y_{5}$) in their terminal-node averages at $x$, inducing dependence through shared weighted outcomes. (B) Partition alignment without overlap: trees are grown on disjoint training sets ($\{1,2,3,4\}$ and $\{6,7,8,9\}$), yet the covariate geometry near $x$ drives both trees to discover the same splits ($X_{1}<c$, then $X_{3}<d$). The prediction point $x$ routes identically through both trees (dashed red paths), landing in structurally equivalent terminal regions. The resulting predictions $T_{\theta}(x)$ and $T_{\theta^{\prime}}(x)$ average different observations drawn from the same neighborhood of $x$, producing dependence through aligned local averaging rules rather than shared outcomes.

### v3

@@ 3.3 Decomposing the within-resample variance term
We refine $V_{\mathrm{in}}(x)$ by conditioning further on the averaging set $A_{\theta}(x)$, which fixes the local grouping used to form the prediction at $x$. Applying the law of total variance with $A_{\theta}(x)$ as the intermediate conditioning object yields
@@
First term: outcome noise through a fixed averaging rule. Once both the resampling pattern and $A_{\theta}(x)$ are fixed, the weights $W_{i}(x;\theta)$ are deterministic and the tree prediction is a fixed linear combination of outcomes. Under conditional independence of outcomes across observational units,
@@
The resampling indicators $\mathcal{I}_{\theta}$ are drawn independently of the outcome-generating process and carry no information about $Y_{i}$ beyond $X$, so they can be removed from the conditioning. The averaging set $A_{\theta}(x)$, by contrast, depends on the realized outcomes through data-adaptive split selection and thus remains. Taking expectations over the design yields
@@
This term measures how much outcome noise at the contributing observations influences the tree prediction through the conditional prediction rule. The conditional variances $\operatorname{Var}(Y_{i}\mid A_{\theta}(x),X)$ are properties of the outcome model at the realized design points; the design controls the weights.
Second term: instability of the grouping rule. Even with the resampling pattern fixed, different realizations of the remaining tree-level randomization can route $x$ into different terminal nodes, inducing different averaging sets $A_{\theta}(x)$. Because $\mathbb{E}[T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x),X]=\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i}$ with deterministic weights, this component measures variability in the prediction at $x$ induced by randomness in which terminal node $x$ is routed to:
@@
This term is zero if and only if the averaging set $A_{\theta}(x)$ is almost surely constant given $(\mathcal{I}_{\theta},X)$, and reflects instability in which observations the tree-growing procedure groups with $x$.
@@ 3.4 Decomposing the resampling component
We next interpret $V_{\mathrm{out}}(x)$, which captures variability induced by the resampling design across trees. Conditioning on the resampling indicators and using the weighted representation,
@@
where $\bar{W}_{i}(x;\mathcal{I}_{\theta})$ is the expected weight assigned to observation $i$ at $x$ under the remaining tree-level randomization, given the resampling realization. Therefore
@@
Different resampling realizations $\mathcal{I}_{\theta}$ change which observations are exposed to the tree, altering the distribution of $A_{\theta}(x)$ and shifting the expected local prediction rule $\bar{W}(x;\mathcal{I}_{\theta})$. Unlike $V_{\mathrm{in}}(x)$, which reflects instability in how a given resampled dataset is locally partitioned, $V_{\mathrm{out}}(x)$ reflects variability from changes in which observations are eligible to contribute to the prediction at $x$.
@@ 4.1 Law of Total Covariance Decomposition
We now decompose the covariance term $C_{T}(x)$ appearing in Theorem 1. From the finite-sample variance identity, $C_{T}(x)$ is the limiting contribution to predictive variability as $B\to\infty$, representing dependence between tree-level conditional predictors that persists under infinite aggregation. Although trees indexed by $\theta$ and $\theta^{\prime}$ are generated independently, their predictions at $x$ are generally dependent because both act on the same realized outcomes under the fixed covariate configuration $X$. To isolate the sources of this dependence, we again condition on the resampling indicators $\mathcal{I}_{\theta}$ and $\mathcal{I}_{\theta^{\prime}}$. Applying the law of total covariance conditional on $X$ with respect to the resampling indicators yields
@@
The two terms in (3) correspond to distinct design mechanisms. The first captures dependence arising from joint inclusion of observations in the averaging sets $A_{\theta}(x)$ and $A_{\theta^{\prime}}(x)$, whereby the same outcomes receive positive weight in both trees. The second captures dependence arising from alignment of the induced conditional predictors at $x$: even when two trees are trained on disjoint subsets of observations, they may group $x$ with observations from the same part of the covariate space, because the underlying signal structure leads both trees to make similar partitioning decisions in the region containing $x$. Resampling schemes govern the first term through joint inclusion probabilities, while feature-level randomization and split selection govern the second through how frequently trees arrive at similar groupings of $x$.
@@ 4.2 Covariance Induced by Shared Training Observations
We first analyze the leading term in (3). This component is nonzero whenever the same observation receives positive weight in both trees at $x$. Joint inclusion in the resampling realizations is necessary but not sufficient; the observation must also lie in the terminal region containing $x$ in both trees. Using the weighted representation $T_{\theta}(x)=\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i}$ and conditioning on $(\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X)$, the covariance expands as
$\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\;\middle|\;\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X\right)$ $=\sum_{i=1}^{n}\sum_{j=1}^{n}W_{i}(x;\theta)\,W_{j}(x;\theta^{\prime})\,\operatorname{Cov}(Y_{i},Y_{j}\mid\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X).$
Because the resampling indicators are drawn independently of the outcome-generating process, conditioning on $(\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}})$ does not impact the joint distribution of $(Y_{i},Y_{j})$ given $X$. Under conditional independence of outcomes across observational units, $\operatorname{Cov}(Y_{i},Y_{j}\mid X)=0$ for $i\neq j$, so the double sum reduces to
@@
where $\sigma_{i}^{2}:=\operatorname{Var}(Y_{i}\mid X)$ is the conditional outcome variance at the realized design point $X_{i}$. Taking expectations over the tree-generating randomization conditional on $X$ yields
@@
The joint weight factor decomposes as
@@
separating joint inclusion under the resampling scheme from the conditional allocation of weight induced by the local partitions. This contribution reflects dependence induced purely by repeated reuse of the same observations in the predictions at $x$, and is identically zero only when the resampling design assigns zero probability to joint inclusion of any observation across trees.
@@ 4.3 Covariance Induced by Partition Alignment
We now analyze the second term in (3), which captures dependence arising from alignment of the induced conditional predictors across independently generated trees. This mechanism does not require any observation to receive positive weight in both trees at $x$. Instead, it arises when distinct trees, potentially trained on disjoint subsets, lead to a similar terminal region containing $x$, therefore yielding similar conditional prediction rules. Two trees trained on different observations may split on the same variables at similar thresholds along the path to $x$, routing $x$ into structurally similar terminal regions; each tree then averages outcomes from the same covariate-defined subpopulation even if the contributing observations are disjoint.
To formalize this, let $R_{\theta}(x)\subset\mathcal{X}$ denote the terminal region containing $x$ under tree draw $\theta$. The averaging set can be written equivalently as
@@
When independent trees induce similar regions $R_{\theta}(x)$ and $R_{\theta^{\prime}}(x)$, they average different observations drawn from essentially the same local subpopulation, producing dependence through aligned partition structure rather than shared outcomes. Define the resampling-conditional prediction
$Z_{\theta}(x):=\mathbb{E}[T_{\theta}(x)\mid\mathcal{I}_{\theta},X]=\sum_{i=1}^{n}\bar{W}_{i}(x;\mathcal{I}_{\theta})\,Y_{i},$
where $\bar{W}_{i}(x;\mathcal{I}_{\theta}):=\mathbb{E}[W_{i}(x;\theta)\mid\mathcal{I}_{\theta},X]$ averages over the distribution of terminal regions induced by split-level randomization, assigning observation $i$ mass proportional to how frequently it is grouped with $x$ across possible tree realizations given the resampling pattern. This is an intermediate quantity between the realized tree-level weight $W_{i}(x;\theta)$ and the forest-level expected weight $\mathbb{E}[W_{i}(x;\theta)\mid X]$, and its introduction makes the alignment mechanism formally identifiable as a distinct component of the covariance structure.
The alignment component is therefore
$C_{\mathrm{align}}(x):=\operatorname{Cov}\!\left(Z_{\theta}(x),Z_{\theta^{\prime}}(x)\;\middle|\;X\right).$
Under conditional independence of outcomes across observational units given $X$,
@@
where $\sigma_{i}^{2}=\operatorname{Var}(Y_{i}\mid X)$. The joint expectation $\mathbb{E}[\bar{W}_{i}(x;\mathcal{I}_{\theta})\,\bar{W}_{i}(x;\mathcal{I}_{\theta^{\prime}})\mid X]$ measures whether independent trees tend to assign predictive weight at $x$ to the same observations after averaging over split-level randomization. When routing of $x$ is stable across tree realizations—so that similar split conditions recur and $x$ repeatedly falls into similar terminal regions—the observations receiving weight, though potentially disjoint across trees, are drawn from the same conditional subpopulation defined by the shared partition structure. Both tree predictions therefore estimate the same conditional mean, yielding positive covariance even without shared training observations.
Alignment reflects a more fundamental source of dependence than observation overlap. Two forests trained on independent samples from the same data-generating process $(Y\mid X)$ will discover similar partitions at $x$ whenever the signal structure produces stable split decisions, because the data-generating process—not the specific training observations—determines which splits are favorable. Observation overlap amplifies this dependence by additionally correlating the noise component: when the same $Y_{i}$ contributes to both predictions at $x$, outcome-level noise does not cancel across trees. Alignment is the primary structural mechanism, present whenever trees are trained on data from the same population; observation overlap provides a secondary reinforcement that requires shared training observations.
@@
Figure 2: Two distinct design-induced dependence mechanisms at a fixed prediction point $x$. (A) Observation overlap: independently generated trees reuse the same outcome (here, $Y_{5}$) in their terminal-node averages at $x$, inducing dependence through shared weighted outcomes. (B) Partition alignment without overlap: trees are grown on disjoint training sets ($\{1,2,3,4\}$ and $\{6,7,8,9\}$), yet the covariate structure at $x$ drives both trees to discover the same splits ($X_{1}<c$, then $X_{3}<d$). The prediction point $x$ routes identically through both trees (dashed red paths), landing in structurally equivalent terminal regions. The resulting predictions $T_{\theta}(x)$ and $T_{\theta^{\prime}}(x)$ average different observations from the same covariate-defined subpopulation at $x$, producing dependence through aligned conditional prediction rules rather than shared outcomes.